OpenStudy (anonymous):

How do you use the unit circle to find the inverse function value in degrees? sin–1 (squareroot of 3 / 2 )

OpenStudy (anonymous):

@phi

OpenStudy (phi):

the unit circle can help you find the correct quadrants, but you have to memorize the actual degrees. You should know the sin, cos and tan of 0º, 30º, 45º, 60º, and 90º there is a pattern for sin: 0 1/2 sqrt(2)/2 sqrt(3)/2 1 if we use sqrt(0) is 0 and sqrt(1) is 1 and the sqrt(4) is 2, the pattern is sqrt(0)/2 sqrt(1)/2 sqrt(2)/2 sqrt(3)/2 sqrt(4)/2 cosine goes in the other direction tangent is sin/cos

OpenStudy (phi):

Can you see what angle has a sin = sqrt(3)/2 ?

OpenStudy (anonymous):

Thats what I cannot understand. No matter how many times I go back in my textbook I cannot wrap my head around any of the trig homework. Its just absolutely confusing me.

OpenStudy (phi):

can you ask a specific question?

OpenStudy (phi):

what is the sin(60º) ? the answer is : memorize it. See above. it is a number between -1 and 1. In this case 0.866025403... and on and on rather than type all the numbers people give the equivalent sqrt(3)/2 type sin(60 degrees)= in the google search window. It will give you that number type sqrt(3)/2= in the google search window. It will give the same number

OpenStudy (phi):

This might help http://www.khanacademy.org/math/trigonometry/v/basic-trigonometry

OpenStudy (phi):

the answer they want (I am assuming) is \[ sin^{-1}\left(\frac{\sqrt{3}}{2} \right)= 60º \text{ and }120º\]

OpenStudy (anonymous):

Will the video better explain how to solve these problems?! Thank you so much. I kept forgetting what website to go to and I've just been having a terrible week

OpenStudy (phi):

I would watch the first two (and maybe the third) they are all short. then look for unit circle on the play list http://www.khanacademy.org/math/trigonometry/v/unit-circle-definition-of-trig-functions-1 is the first one. Go get some popcorn and see how it goes...

OpenStudy (anonymous):

Thank you so much! I'll watch them right away! Also one of my questions I'm supposed to use a calculator find the values of the inverse function in radians, the calculator only gives you one answer...it asks for two. How would I figure out the second one?

OpenStudy (phi):

The second answer is in a different quadrant, and depends on what trig function you are doing

OpenStudy (anonymous):